# Near Exact Privacy Amplification for Matrix Mechanisms

### v1

@@ Abstract
We study the problem of computing the privacy parameters for DP machine learning when using privacy amplification via random batching and noise correlated across rounds via a correlation matrix $\mathbf{C}$ (i.e., the matrix mechanism). Past work on this problem either only applied to banded $\mathbf{C}$, or gave loose privacy parameters. In this work, we give a framework for computing near-exact privacy parameters for any lower-triangular, non-negative $\mathbf{C}$. Our framework allows us to optimize the correlation matrix $\mathbf{C}$ while accounting for amplification, whereas past work could not. Empirically, we show this lets us achieve smaller RMSE on prefix sums than the previous state-of-the-art (SOTA). We also show that we can improve on the SOTA performance on deep learning tasks. Our two main technical tools are (i) using Monte Carlo accounting to bypass composition, which was the main technical challenge for past work, and (ii) a “balls-in-bins” batching scheme that enables easy privacy analysis and is closer to practical random batching than Poisson sampling.
@@ 1.1.1 Algorithmic Contributions
Near-exact analysis via Monte Carlo: We recognize a common issue with prior amplification analysis of correlated noise mechanisms, which is their reliance on composition which led to slack for various reasons (see Section 1.2). We thus show how to use Monte Carlo accounting (Wang et al., 2023) which instead observes that approximate DP guarantees can be written as the expected value of some function of the privacy loss and $\varepsilon$ (see Section 2 for more details). Because it is easy to do Monte Carlo accounting for the whole matrix $\mathbf{C}\mathbf{x}+\mathbf{z}$, we can bypass the need for composition. In turn, we are able to obtain near-exact privacy analysis (only paying for the sampling error of Monte Carlo approximation) while supporting general $\mathbf{C}$, improving upon the weaknesses of both the past works.
@@ 2 Privacy Loss Distributions and Monte Carlo Accounting
A mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if for any adjacent databases $D\sim D^{\prime}$, we have $H_{e^{\varepsilon}}(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\leq\delta$. Suppose $P,Q$ are a dominating pair (Zhu et al., 2022) for $\mathcal{M}$ of interest; that is, for any $D\sim D^{\prime},\alpha\geq 0$ we have $H_{\alpha}(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\leq H_{\alpha}(P,Q)$. Then, by sampling $X\sim P$ and computing $Y=\log P(X)/Q(X)$, we then say that the privacy-loss distribution (PLD) (Balle et al., 2018) of $P$ and $Q$ is the law of $Y$.
Now to prove $(\varepsilon,\delta)$-DP for $\mathcal{M}$, it suffices to show $H_{e^{\varepsilon}}(P,Q)\leq\delta$. The main observation in Monte Carlo accounting is that we can write $H_{\alpha}$ as an expectation over $Y$,
$H_{\alpha}(P,Q)=\mathbb{E}_{Y}[\max\{1-\alpha e^{-Y},0\}].$
Then, if exactly computing $H_{\alpha}(P,Q)$ is hard but sampling $Y$ is easy, we can estimate $\delta$ for some $\varepsilon$ by taking the average of sufficiently many samples of $\max\{1-e^{\varepsilon-Y},0\}$. Namely, let $\hat{\delta}$ be the estimated $\delta$ value, i.e. the average over some samples of $\max\{1-e^{\varepsilon-Y},0\}$. As a starting point, we can already claim $(\varepsilon,\hat{\delta})$-DP as an informal privacy statement. To make it formal, we can use the Estimate-Verify-Release framework of (Wang et al., 2023) (Algorithm 1): we fix a target $(\varepsilon,\delta)$-DP privacy guarantee ahead of time, and use Monte Carlo accounting to verify that $\mathcal{M}$ satisfies with $(\varepsilon,\delta)$-DP, and only run $\mathcal{M}$ if the verification succeeds.
@@ 3 Balls-in-bins batching
Figure 1: A comparison of the how different amplification methods might form batches. Here, we use $n=6$ rounds, and form $b=3$ batches per epoch for $E=2$ epochs. Note that shuffling uses fixed batch sizes, and all methods but Poisson sampling use the same batching across epochs and enforce exactly one participation per example per epoch.
@@
Suppose $\mathbf{C}\in\mathbb{R}^{n\times n}$ is a lower-triangular matrix with non-negative entries. Then $P,Q$ as defined above is a dominating pair for the correlated noise mechanism with balls-in-bins batching.
@@
Because each user is assigned to their batch independently, we can assume without loss of generality that contributions from all users other than the differing user is always 00 (for a justification of this reduction, see e.g. the first part of proof of Theorem 4.8 in Choquette-Choo et al. (2024b)). Then, for the input matrix $\mathbf{x}\in\mathbb{R}^{n\times p}$, the rows have at most unit norm in the entries where the differing user participates and 0 everywhere else. If the differing user is assigned to the $i$th batch, it is immediate by triangle inequality that ${\left\|{(\mathbf{C}\mathbf{x})_{k,:}}\right\|}\leq\sum_{j=0}^{E-1}|\mathbf{C}|_{k,b\cdot j+i}$. Because every user participates in each batch with probability $\frac{1}{b}$, we can apply Lemma 4.5 of Choquette-Choo et al. (2024b) to get that the correlated noise mechanism with balls-with-bins batching is dominated by the MoG mechanism (using their notation) $\mathcal{M}_{MoG}(\{\frac{1}{b},\dots,\frac{1}{b}\},\{\sum_{j=0}^{E-1}|\mathbf{C}|_{1:n,b\cdot j+i}\}_{1\leq i\leq b})$. This MoG is exactly $P$, so we have the desired result. ∎
@@ 4.1 Calibrating $\sigma$
Hence we can sample $Y\sim PLD(P_{\mathbf{C},\sigma},Q_{\sigma})$ by first sampling $X\sim P_{\mathbf{C},\sigma}$ and computing $Y=\log\left(\frac{P_{\mathbf{C},\sigma}}{Q_{\sigma}}(X)\right)$. Note that computing $Y$ is efficient since $P_{\mathbf{C},\sigma}$ has only $b$ modes. To make the sampling independent of $\sigma$, we can instead sample $i\sim Uni([b])$ and $Z\sim N(0,I)$ then for any $\sigma$ compute $Y=\log\left(\frac{P_{\mathbf{C},\sigma}}{Q_{\sigma}}(\mathbf{m}_{i}+\sigma\cdot Z)\right)$. This process is summarized in Algorithm 2, which uses the function $\hat{\delta}$ to estimate the $\delta$ at a fixed $\sigma$. The inner computation of $\hat{\delta}$ is easily parallelizable, as it is a simple average over a given sample size. On top of this, a bisection algorithm is run, which only requires $O(\log(1/\delta))$ many iterations typically because $\sigma$ is a scalar.
@@ 5.1.1 Analysis of the amplification schemes
(b) Improvement from optimizing $\mathbf{C}$ using objective accounting for amplification.
@@ 5.1.2 Improvements due to optimization
We now look at the benefits of using our optimization procedure to choose $\mathbf{C}$. We compare two variants of our optimization scheme, one which optimizes over general Toeplitz matrices, and another which optimizes over the BLT matrix family defined by Dvijotham et al. (2024). We use the 3-buffer variant of BLTs, i.e. they can be specified using 6 parameters and only require an overhead of 3 times the model size to store the state required for correlated noise generation. In Figure 3(b) we plot the improvement in RMSE for both these variants relative to the best $b$-banded $\mathbf{C}$ amplified using the sampling scheme and analysis of Choquette-Choo et al. (2024a).
@@ 5.2 CIFAR Results
For balls-in-bins batching, we conservatively set the target $\delta$ in accounting to be $8\cdot 10^{-6}$, and use sufficiently many samples in Monte Carlo estimation such that the wrapper of Wang et al. (2023) gives the desired $(\varepsilon,10^{-5})$-DP guarantee and has a small failure probability.

### v2

@@ Abstract
We study the problem of computing the privacy parameters for DP machine learning when using privacy amplification via random batching and noise correlated across rounds via a correlation matrix $\mathbf{C}$ (i.e., the matrix mechanism). Past work on this problem either only applied to banded $\mathbf{C}$, or gave loose privacy parameters. In this work, we give a framework for computing near-exact privacy parameters for any lower-triangular, non-negative $\mathbf{C}$. Our framework allows us to optimize the correlation matrix $\mathbf{C}$ while accounting for amplification, whereas past work could not. Empirically, we show this lets us achieve smaller RMSE on prefix sums than the previous state-of-the-art (SOTA). We also show that we can improve on the SOTA performance on deep learning tasks. Our two main technical tools are (i) using Monte Carlo accounting to bypass composition, which was the main technical challenge for past work, and (ii) a “balls-in-bins” batching scheme that enables easy privacy analysis and implementation-wise is closer to shuffling (widely considered a practical random batching method) than Poisson sampling.
@@ 1.1.1 Algorithmic Contributions
Near-exact analysis via Monte Carlo: We recognize a common issue with prior amplification analysis of correlated noise mechanisms, which is their reliance on composition which led to slack for various reasons (see Section 1.2). We thus show how to use Monte Carlo accounting (Wang et al., 2023) which instead observes that approximate DP guarantees can be written as the expected value of some function of the privacy loss and $\varepsilon$ (see Section 2 for more details). Because it is easy to do Monte Carlo accounting for the whole matrix $\mathbf{C}\mathbf{x}+\mathbf{z}$, we can bypass the need for composition. In turn, we are able to obtain near-exact privacy analysis (only paying for the sampling error of Monte Carlo approximation, which can be made arbitrarily small) while supporting general $\mathbf{C}$, improving upon the weaknesses of both the past works.
@@ 1.2 Background and Prior Work
1.3 Comparison to Chua et al.
An independent and concurrent work of Chua et al. (2025) also proposed balls-in-bins sampling, defined identically to our work. We believe the fact that both papers independently arrived at the same batch sampling method is evidence that it is a natural method to consider. Similarly to our work, Chua et al. (2025) use Monte Carlo-based privacy accounting to handle balls-in-bins sampling. In their work, they consider the independent noise setting ($\mathbf{C}=\mathbf{I}$, as opposed to general $\mathbf{C}$ in our work), which enables them to make their sampler more efficient by (i) using an importance sampler that reduces the variance of the Monte Carlo accountant, and (ii) an “order statistics sampling” method which estimates a pessimistic value of $\delta$ but is more efficient than the unbiased sampler. To the best of our knowledge, their techniques do not easily translate to the setting of general $\mathbf{C}$. It is an interesting question to see if the benefits of both works can be combined, i.e., if we can design more efficient samplers for Monte Carlo accounting in the setting of general $\mathbf{C}$.
@@ 2 Privacy Loss Distributions and Monte Carlo Accounting
A mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if for any adjacent databases $D\sim D^{\prime}$ (under the add-remove adjacency), we have $H_{e^{\varepsilon}}(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\leq\delta$. Suppose $P,Q$ are a dominating pair (Zhu et al., 2022) for $\mathcal{M}$ of interest; that is, for any $D\sim D^{\prime},\alpha\geq 0$ we have $H_{\alpha}(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\leq H_{\alpha}(P,Q)$. Then, by sampling $X\sim P$ and computing $Y=\log P(X)/Q(X)$, we then say that the privacy-loss distribution (PLD) (Balle et al., 2018) of $P$ and $Q$ is the law of $Y$.
Now to prove $(\varepsilon,\delta)$-DP for $\mathcal{M}$, it suffices to show $H_{e^{\varepsilon}}(P,Q)\leq\delta$. The main observation in Monte Carlo accounting is that we can write $H_{\alpha}$ as an expectation over $Y$, $H_{\alpha}(P,Q)=\mathbb{E}_{Y}[\max\{1-\alpha e^{-Y},0\}]$. Then, if exactly computing $H_{\alpha}(P,Q)$ is hard but sampling $Y$ is easy, we can estimate $\delta$ for some $\varepsilon$ by taking the average of sufficiently many samples of $\max\{1-e^{\varepsilon-Y},0\}$. Namely, let $\hat{\delta}$ be the estimated $\delta$ value, i.e. the average over some samples of $\max\{1-e^{\varepsilon-Y},0\}$. As a starting point, we can already claim $(\varepsilon,\hat{\delta})$-DP as an informal privacy statement. To make it formal, we can use the Estimate-Verify-Release framework of (Wang et al., 2023) (Algorithm 1): we fix a target $(\varepsilon,\delta)$-DP privacy guarantee ahead of time, and use Monte Carlo accounting to verify that $\mathcal{M}$ satisfies with $(\varepsilon,\delta)$-DP, and only run $\mathcal{M}$ if the verification succeeds.
@@
The following theorem can be used to derive the tail bound required for Theorem 2.1:
Theorem 2.2.
Suppose we use $s$ samples in computing $\hat{\delta}$ in Algorithm 1. Then for any $P,Q$ such that $H_{\varepsilon}(P,Q)\geq\tau\delta,\tau>1$, we have $\mathop{\mathbf{Pr}}[\hat{\delta}<\delta]\leq\exp\left(-\frac{s(\tau-1)^{2}\delta}{8\tau/3-2/3}\right).$
Proof.
If $H_{\varepsilon}(P,Q)>\tau\delta$, then we can express $\hat{\delta}=\frac{1}{s}\sum_{i=1}^{s}\hat{\delta}_{i}$, where each $\hat{\delta}_{i}$ is an i.i.d. random variable in the range $[0,1]$ with mean $\mu\geq\tau\delta$. By the Bhatia-Davis inequality, we have $\mathbf{Var}\left(\hat{\delta}_{i}\right)\leq(1-\mu)\mu$. So $\mathbb{E}[\hat{\delta}_{i}^{2}]=\mathbb{E}[\hat{\delta}_{i}^{2}]^{2}+\mathbf{Var}\left(\hat{\delta}_{i}\right)=\mu^{2}+\mathbf{Var}\left(\hat{\delta}_{i}\right)\leq\mu$.
Now we can apply Bernstein’s inequality to get that:
$\mathop{\mathbf{Pr}}[\hat{\delta}<\mu-t]\leq\exp\left(-\frac{st^{2}/2}{\mathbb{E}[\hat{\delta}_{1}^{2}]+t/3}\right)\leq\exp\left(-\frac{st^{2}/2}{\mu+t/3}\right).$
Setting $t=\mu-\delta$:
$\mathop{\mathbf{Pr}}[\hat{\delta}<\delta]\leq\exp\left(-\frac{s(\mu-\delta)^{2}/2}{4\mu/3-\delta/3}\right).$
Since $\mu\geq\tau\delta>\delta$, we have:
$\frac{d}{d\mu}\frac{s(\mu-\delta)^{2}/2}{4\mu/3-\delta/3}=\frac{3s(\mu-\delta)(2\mu-\delta)}{(4\mu-3\delta)^{2}}>0,$
i.e. the bound $\mathop{\mathbf{Pr}}[\hat{\delta}<\delta]$ is decreasing in $\mu$, so it is minimized by setting $\mu=\tau\delta$. Plugging this value of $\mu$ in gives as desired:
$\mathop{\mathbf{Pr}}[\hat{\delta}<\delta]\leq\exp\left(-\frac{s(\tau-1)^{2}\delta}{8\tau/3-2/3}\right).$
@@ 3 Balls-in-bins batching
Figure 1: A comparison of the how different amplification methods might form batches. Here, we use $n=6$ rounds, and form $b=3$ batches per epoch for $E=2$ epochs (visually represented as one row for each epoch). Note that shuffling uses fixed batch sizes, and all methods but Poisson sampling use the same batching across epochs and enforce exactly one participation per example per epoch.
@@
Suppose $\mathbf{C}\in\mathbb{R}^{n\times n}$ is a lower-triangular matrix with non-negative entries. Then $P,Q$ (resp. $Q,P$) as defined above is a dominating pair for the correlated noise mechanism with balls-in-bins batching under the add (resp. remove) adjacency.
@@
We will use Lemma 4.5 of Choquette-Choo et al. (2024b), restated here for convenience:
Lemma 3.3.
Let $\mathbf{c}_{1},\ldots,\mathbf{c}_{k}\in\mathbb{R}^{n\times p}$. Let $\mathbf{c}_{1}^{\prime},\ldots,\mathbf{c}_{k}^{\prime}\in\mathbb{R}^{n}$ be such that $\left\|\mathbf{c}_{i}[j,:]\right\|_{2}\leq\mathbf{c}_{i}^{\prime}(j)$ for all $i,j$. Then letting
$P=N(0,\sigma^{2}\mathbb{I}_{(n\times p)\times(n\times p)}),Q=\sum_{i}p_{i}N(\mathbf{c}_{i},\sigma^{2}\mathbb{I}_{(n\times p)\times(n\times p)})$
$P^{\prime}=N(0,\sigma^{2}\mathbb{I}_{n\times n}),Q^{\prime}=\sum_{i}p_{i}N(\mathbf{c}_{i}^{\prime},\sigma^{2}\mathbb{I}_{n\times n}),$
for all $\alpha$ we have $H_{\alpha}(P,Q)\leq H_{\alpha}(P^{\prime},Q^{\prime})$. Furthermore, this holds even if the $j$th row of each $\mathbf{c}_{i}$ is chosen as a function of the first $j-1$ rows of $P,Q$ (subject to $\left\|\mathbf{c}_{i}[j,:]\right\|_{2}\leq\mathbf{c}_{i}^{\prime}(j)$) while $\mathbf{c}_{i}^{\prime}$ remain fixed.
We give the proof for the add adjacency. Proving $P,Q$ is a dominating pair for the add adjacency implies $Q,P$ is a dominating pair for the remove adjacency by Lemma 29 of Zhu et al. (2022).
Because each user is assigned to their batch independently, we can assume without loss of generality that contributions from all users other than the differing user are always 00. In more detail, by post-processing, we can assume that we release the contributions to the input matrix of all examples except the differing user’s. Let $\mathbf{x}$ be these contributions, and $\mathbf{x}^{\prime}$ be $\mathbf{x}$ plus the contribution of the differing user. Then distinguishing $\mathbf{C}\mathbf{x}+\mathbf{z}$ and $\mathbf{C}\mathbf{x}^{\prime}+\mathbf{z}$ is equivalent to distinguishing $\mathbf{z}$ and $\mathbf{C}(\mathbf{x}^{\prime}-\mathbf{x})+\mathbf{z}$, exactly the setting where all users except the differing user only contribute 00.
Now, for the input matrix $\mathbf{x}\in\mathbb{R}^{n\times p}$, the rows have at most unit norm in the entries where the differing user participates and 0 everywhere else. If the differing user is assigned to the $i$th batch, it is immediate by triangle inequality that ${\left\|{(\mathbf{C}\mathbf{x})_{k,:}}\right\|}\leq\sum_{j=0}^{E-1}|\mathbf{C}|_{k,b\cdot j+i}$. Because every user participates in each batch with probability $\frac{1}{b}$, we can apply Lemma 4.5 of Choquette-Choo et al. (2024b) to exactly get that the correlated noise mechanism with balls-with-bins batching is dominated by the pair $P,Q$. ∎
For simplicity of presentation, we will focus on the add adjacency in the rest of the paper. It is easy to extend results to the remove adjacency, and for the add-and-remove adjacency we can simply take the worse of the two adjacencies’ privacy guarantees (paying a factor of 2 in the failure probability of Theorem 2.2 by a union bound). We also remark that 3.2 (and any results in our paper for the add/remove adjacency) can be readily extended to the replace adjacency using recent results of Schuchardt et al. (2024).
@@ 4.1 Calibrating $\sigma$
Hence we can sample $Y\sim PLD(P_{\mathbf{C},\sigma},Q_{\sigma})$ by first sampling $X\sim P_{\mathbf{C},\sigma}$ and computing $Y=\log\left(\frac{P_{\mathbf{C},\sigma}}{Q_{\sigma}}(X)\right)$. Note that computing $Y$ is efficient since $P_{\mathbf{C},\sigma}$ has only $b$ modes. To make the sampling independent of $\sigma$, we can instead sample $i\sim Uni([b])$ and $Z\sim N(0,I)$ then for any $\sigma$ compute $Y=\log\left(\frac{P_{\mathbf{C},\sigma}}{Q_{\sigma}}(\mathbf{m}_{i}+\sigma\cdot Z)\right)$. This process is summarized in Algorithm 2, which uses the function $\hat{\delta}$ to estimate the $\delta$ at a fixed $\sigma$. The inner computation of $\hat{\delta}$ is easily parallelizable, as it is a simple average over a given sample size. On top of this, a bisection algorithm is run, which can converge [^3] To show bisection search converges, if $\sigma_{max}$ is the $\sigma$ achieving the target DP guarantee without amplification (which we easily can compute using existing accounting libraries), $\hat{\delta}(0;\ldots)=1>\delta$ and $\hat{\delta}(\sigma_{max};\ldots)<\delta$ (whp). In turn, $\hat{\delta}(\sigma;\ldots)=\delta$ for some $\sigma\in[0,\sigma_{max}]$ and bisection converges to this point by continuity of $\hat{\delta}$. to within $\Delta$ additive error in only $O(\log(1/\Delta))$ many iterations because $\sigma$ is a scalar.
@@ 5 Experiments
For all our empirical results, we use $\delta=8\cdot 10^{-6},\tau=1.25,s=10^{8}$. For this set of values, Theorem 2.2 gives a failure probability less than $7.2\cdot 10^{-9}$. We use Algorithm 1 to verify the DP guarantee for both the add and remove adjacencies, but by a union bound the failure probability increases to at most $1.5\cdot 10^{-8}$ once accounting for both adjacencies, much smaller than $\tau\delta=10^{-5}$. So, we can apply Theorem 2.1 to show $(\varepsilon,10^{-5})$-DP for all our empirical results.
@@ 5.1.1 Analysis of the amplification schemes
In Figure 3(c) we repeat this comparison but fixing $\varepsilon$ and varying the sampling probability / iterations per epoch instead. We fix 128 rounds of training, and fix a target DP guarantee of $(1,10^{-5})$-DP. We vary the sampling probability $p$ and compare the RMSE achieved by the best banded matrix with either balls-in-bins sampling + our amplification analysis or the sampling scheme and analysis of Choquette-Choo et al. (2024a). For balls-in-bins, we treat $1/b$ where $b$ is the iterations per epoch as the sampling probability for the purpose of plotting, i.e. at each point on the x-axis both algorithms have the same expected batch size.
@@
(b) For each of the amplification methods in Figure 3(a), the best choice of $b$ at each value of $\varepsilon$.
(c) Comparison of improvement due to amplification as a function of sampling probability.
(d) Improvement from optimizing $\mathbf{C}$ using objective accounting for amplification.
@@ 5.1.2 Improvements due to optimization
We now look at the benefits of using our optimization procedure to choose $\mathbf{C}$. We compare two variants of our optimization scheme, one which optimizes over general Toeplitz matrices, and another which optimizes over the BLT matrix family defined by Dvijotham et al. (2024). We use the 3-buffer variant of BLTs, i.e. they can be specified using 6 parameters and only require an overhead of 3 times the model size to store the state required for correlated noise generation. In Figure 3(d) we plot the improvement in RMSE for both these variants relative to the best $b$-banded $\mathbf{C}$ amplified using the sampling scheme and analysis of Choquette-Choo et al. (2024a).
@@ 5.2 CIFAR Results
Method / $\varepsilon$ 0.5 1.0 2.0 4.0 8.0
BandMF + Poisson 1.018 0.778 0.606 0.481 0.388
BandMF + Balls-in-Bins 2.829 2.071 1.455 0.802 0.470
Optimized BLT + Balls-in-Bins 2.609 1.598 1.079 0.694 0.448
Optimized Toeplitz + Balls-in-Bins 2.003 1.182 0.864 0.639 0.445
Figure 7: Noise multiplier $\sigma$ used in the CIFAR10 experiments for different correlated noise strategies and amplification methods. Note that different strategies which have higher noise cancellation require higher $\sigma$, so a lower noise multiplier alone does not imply better learning performance.
Method / $\varepsilon$ 0.5 1.0 2.0 4.0 8.0
BandMF + Poisson 2 4 8 16 32
BandMF + Balls-in-Bins 16 32 64 64 64
Figure 8: Number of bands used in the CIFAR10 experiments for BandMF combined with different amplification methods.